OpenStudy (anonymous):

cos(arcsin(-sqrt2/2)) inverse trig function help

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

show me what you have so far

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

arcsin(-sqrt2/2)= -pi/4. Cos(-pi/4) = sqrt2/2 ??

jimthompson5910 (jim_thompson5910):

your steps and final answer is correct

OpenStudy (anonymous):

niceee

OpenStudy (anonymous):

cot^-1(tan pi/3)

OpenStudy (anonymous):

for this problem i got pi/6. ??

jimthompson5910 (jim_thompson5910):

also correct

OpenStudy (anonymous):

some person on here said it was wrong. are u sure its correct?

jimthompson5910 (jim_thompson5910):

wolfram alpha confirms it http://www.wolframalpha.com/input/?i=arccot%28tan%28%281%2F3%29*Pi%29%29

OpenStudy (anonymous):

cool thanks. can u help me out on arcsin(sin 3pi)

jimthompson5910 (jim_thompson5910):

sin(3pi) = ???

OpenStudy (anonymous):

umm 0/

OpenStudy (anonymous):

0 ?

jimthompson5910 (jim_thompson5910):

yes now compute arcsin(0)

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

i have no idea. im currently looking at my unit circle

jimthompson5910 (jim_thompson5910):

yes so the final answer is 0

jimthompson5910 (jim_thompson5910):

sin(0) = 0 arcsin(0) = 0

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

im working on some others at this moment

OpenStudy (anonymous):

cos(sin^-1 sqrt3/2). i got 1/2??

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

cot(sec^-1 2). i got sqrt3/3??

jimthompson5910 (jim_thompson5910):

you're on a roll

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

arccot(-sqrt3/3). i got pi/3??

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

it was good while it lasted lol

jimthompson5910 (jim_thompson5910):

hint: cotangent is negative in Q2 and Q4 range of arccotangent is `0 < y < pi` so the output of arccotangent will be in Q2 if the input is negative

OpenStudy (anonymous):

so is it - pi/3 ?

jimthompson5910 (jim_thompson5910):

no, the angle will be somewhere in Q2

OpenStudy (anonymous):

5pi/3??

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

2pi/3??

jimthompson5910 (jim_thompson5910):

yep 2pi/3

OpenStudy (anonymous):

how though?

jimthompson5910 (jim_thompson5910):

because cot(2pi/3) = -sqrt(3)/3

OpenStudy (anonymous):

so does 5pi/3

jimthompson5910 (jim_thompson5910):

recall that cot(x) = 1/tan(x) or cot(x) = cos(x)/sin(x)

jimthompson5910 (jim_thompson5910):

but 5pi/3 is in Q4

OpenStudy (anonymous):

i dont understand the importance of the quadrants

OpenStudy (anonymous):

how do i know which one to use?

jimthompson5910 (jim_thompson5910):

the range of arccot is 0 < y < pi, so that covers Q1 and Q2

jimthompson5910 (jim_thompson5910):

whatever the output of arccot it will either be in Q1 or Q2

OpenStudy (anonymous):

will i ever use q3 or q4?

jimthompson5910 (jim_thompson5910):

not with arccot

jimthompson5910 (jim_thompson5910):

read this article https://en.wikipedia.org/wiki/Inverse_trigonometric_functions look at the table that says "The principal inverses are listed in the following table."

OpenStudy (anonymous):

all right. sin^-1(sin 7pi/4). i got pi/4. probably wrong. dont really know this one either.

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

what is sin(7pi/4) equal to

OpenStudy (anonymous):

-sqrt2/2

jimthompson5910 (jim_thompson5910):

now compute arcsin(-sqrt(2)/2)

jimthompson5910 (jim_thompson5910):

range of arcsin: -pi/2 <= y <= pi/2 the input is negative, so the angle is in Q4

OpenStudy (anonymous):

ummm 7pi/4??